# A Fermi Model of Quantum Black Hole

### v2

@@ II.5 Black hole with a cosmological constant
In addition, there is the background pressure (II.54). As for the gravitational energy, we note that the definition (II.14) requires an asymptotic flat background as a reference level for the energy. Now since there is no asymptotically flat region in the de Sitter black hole spacetime, an appropriate modification to the definition of the black hole mass is to subtract out the background contribution from the pure de Sitter space ($M=0$) and use
$E_{g}:=\frac{\kappa A}{4\pi G}-\left(\frac{\kappa A}{4\pi G}\right)_{M=0}.$ (II.60)
@@
$E_{g}=M$ (II.61)
and hence the pressure
$P_{g}=\frac{1}{8\pi GR^{2}}-\frac{\Lambda}{8\pi G}$ (II.62)
@@
$P_{U}+P_{\Lambda}=P_{g}$ (II.63)
and our model for the black hole is stable.
@@
$ds^{2}=-(1-\frac{2GM}{r}+\frac{|\Lambda|\,r^{2}}{3})dt^{2}+(1-\frac{2GM}{r}+\frac{|\Lambda|\,r^{2}}{3})^{-1}dr^{2}+r^{2}d\Omega^{2}.$ (II.64)
@@
$01-\frac{2GM}{R}+\frac{|\Lambda|R^{2}}{3}=$ 0 (II.65)
and the black hole is situated in the background AdS space with a boundary at infinity. We model the black hole interior with a Fermi system of the same properties (II.57), (II.58), (II.59). The gravitational energy is defined the same way (II.60) with a subtraction and we obtain
$P_{g}=\frac{1}{8\pi GR^{2}}+\frac{|\Lambda|}{8\pi G}.$ (II.66)
Pressure is again balanced and the Fermi model of black hole interior is stable against collapse. The Bekenstein-Hawking entropy works out the same way as before.

### v3

@@ II.3 Schwarzschild black hole as a system of Fermi quanta
We remark that the area law receives corrections in higher derivative modified gravity. In principle, such corrections can be accounted for in our model by using a modified Fermi level and a modified density of states. However deriving these results will require the knowledge of a quantum theory of gravity. Please see [32, 33] for a recent proposal of a theory of quantum gravity as a quantum mechanics of quantum spaces. It is quite amusing that most of the features of the phenomenological Fermi model appear there in a consistent setup.
@@ II.5 Black hole with a cosmological constant
In addition, the system has an energy
$E_{\Lambda}=\frac{\Lambda}{6G}R^{3}$ (II.60)
and the pressure (II.54) due to the presence of a cosmological constant. As for the gravitational energy, we note that the definition (II.14) requires an asymptotic flat background as a reference level for the energy. Now since there is no asymptotically flat region in the de Sitter black hole spacetime and also because there is not a timelike Killing vector in the global de Sitter spacetime, the definition of mass in de Sitter spacetime is non-trivial. One can employ a convenient definition of the black hole mass by subtracting away the background contribution from the pure de Sitter space ($M=0$):
$E_{g}:=\frac{\kappa A}{4\pi G}-\left(\frac{\kappa A}{4\pi G}\right)_{M=0}.$ (II.61)
@@
$E_{g}=M.$ (II.62)
Actually one can use any other definition as long as the relation (II.62) is reproduced. The principle of equivalence then states that equality $U=E_{g}+E_{\Lambda}$ between the energy of the Fermi system and the total gravitational energy of the black hole solution. This gives immediately the relation (II.55) satisfied by the horizon radius of a dS black hole. At the same time, we have from (II.62) the gravitational pressure
$P_{g}=\frac{1}{8\pi GR^{2}}-\frac{\Lambda}{8\pi G}$ (II.63)
@@
$P_{U}=P_{g}+(-P_{\Lambda}),$ (II.64)
where the term $-P_{\Lambda}$ is the pressure acting on the system from externally due to the background energy density. Our model for the black hole is stable. We remark that as the Fermi degrees of freedom is supposed to be elementary quanta of the quantized space, it should be possible to build a Fermi model that describes the cosmological horizon as well. This is an interesting problem to consider for the quantum mechanics of quantum space [32, 33].
@@
$ds^{2}=-(1-\frac{2GM}{r}+\frac{|\Lambda|\,r^{2}}{3})dt^{2}+(1-\frac{2GM}{r}+\frac{|\Lambda|\,r^{2}}{3})^{-1}dr^{2}+r^{2}d\Omega^{2}.$ (II.65)
@@
$01-\frac{2GM}{R}+\frac{|\Lambda|R^{2}}{3}=$ 0 (II.66)
and the black hole is situated in the background AdS space with a boundary at infinity. We model the black hole interior with a Fermi system of the same properties (II.57), (II.58), (II.59). The gravitational energy for AdS black hole is well defined and one have
$E_{g}=M.$ (II.67)
Equating the energy of the Fermi system $\frac{R}{2G}$ with the total gravitational energy $M-\frac{|\Lambda|}{6G}R^{3}$ of the black hole spacetime, we obtain the mass-horizon radius relation for AdS black hole. In the same way, we have the gravitational pressure
$P_{g}=\frac{1}{8\pi GR^{2}}+\frac{|\Lambda|}{8\pi G}$ (II.68)
in the outside region of the black hole. Pressure is again balanced and the Fermi model of black hole interior is stable against collapse. The Bekenstein-Hawking entropy works out the same way as before.